# Singleton Park, Swansea SA2 8PP, UK

### v1

@@ 2.3 Precise Formulation of Scattering Amplitudes
In the strict high-energy fixed-angle limit $y_{0}\to\infty$, and our predictions do not depend on the exact value of $\tilde{\alpha}^{\prime}$, as long as it remains finite. Hence, the leading behavior does not depend on the value of the IR cut-off, or more generally, on how confinement is realized in the model.
When examining pion scattering, we must also account for the pion’s isospin (realized in our holographic model using Chan–Paton factors, which we omitted earlier in Section 2.2). To do this, we follow the standard procedure Ecker:1994gg; Igi:1998gn, in which amplitudes of various pion scattering processes are represented by a single analytic function, the invariant amplitude:
@@ 2.3.1 Regularisation and Numerical Methods
As mentioned, evaluating equation (36) at integer values of $\tilde{\alpha}^{\prime}s$ leads to singularities that cannot be regularised.  These can be shown to correspond to logarithmic $s$-channel singularities Bianchi:2021sug, although we do not expect that these singularities are physical, as they are restricted to finite $s$.  On the other hand, as long as $s$ does not coincide with a singularity, i.e.  when $\tilde{\alpha}^{\prime}s\neq n$, all other singularities can be regularised to give a well-defined integral, from which the angular dependence of the amplitude can be computed.  One way to achieve this is by shifting all the poles by $i\epsilon$, effectively giving a small width to all the mesons on the Regge trajectory. [^5] Phenomenologically, we expect finite widths, since all of the states on the $\rho$-meson Regge trajectory are unstable and decay.  In fact, the string decay width is expected to increase with spin, being proportional to the string length, which scales linearly with the mass Dai:1989cp; Peeters:2005fq; Iengo:2006gm; Sonnenschein:2017ylo.  In the ’t Hooft limit, the excited states are stable Witten:1979kh, since the width is $\sim 1/N$, and the width functions serve only as regulators, which are eventually taken to zero.  A finite physical limit can then be defined by taking the limit $\epsilon\to 0$ in an appropriate manner.
@@
Now, since our predictions are based on tree-level superstring scattering amplitudes, the pion scattering amplitude $\mathcal{A}$ must be real.  This follows from the unitarity constraint Peskin:1995ev:
@@
where $T$ is the transfer matrix related to $\mathcal{A}$ via the LSZ formula, so $\mathcal{A}\sim T$.  At tree level, the right-hand side above vanishes, implying that unitarity forces $\mathcal{A}$ to be real. This connection between unitarity and the reality of the amplitude can also be seen from the fact that, at large-$N$, the poles of the amplitude must lie on the positive real axis Caron-Huot:2016icg.
@@ 3 Background Material on Experimental Data
In this section, we describe how the data for $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$ scattering can be extracted, under certain approximations, from the process $\pi^{-}p\to\pi^{+}\pi^{-}n$ studied in reference Bromberg:1983he. To this end, we introduce two models of increasing generality for extracting the data that, together with our holographic model, will form the basis for the analysis in Section 4.
In the experiment reported in reference Bromberg:1983he, the process $\pi^{-}p\to\pi^{+}\pi^{-}n$ was studied at incident beam momenta $|\vec{q}_{\text{l.f.}}|$ of approximately 100 GeV and 175 GeV in the lab frame. The incident beam was primarily composed of pions (approximately 90% at 100 GeV and 92% at 175 GeV), while liquid hydrogen was used as the target.  All the reported data were corrected for known experimental biases and acceptance losses, with uncertainties in these corrections remaining significantly smaller than the statistical errors across all data sets. Following the conventions of reference Bromberg:1983he, we frequently use the Gottfried–Jackson frame in this section (see Figure 2).  Additional information about the experiment can be found in references Fredericksen:1982hb; Stampke:1982cf; Bromberg:1983kv; Bromberg:1984ny, while related studies are reported in references Aderholz:1964tpc; Aderholz:1964lgn; Bondar:1966zz; Hyams:1968zza; Robertson:1973tk; Grayer:1974cr; Birmingham-Rutherford-TelAviv-Westfield:1977ohi; Wicklund:1977in; Weilhammer:1979wg.
@@
After averaging over the target proton spin and summing over the recoil neutron spin, the process $\pi^{-}p\to\pi^{+}\pi^{-}n$ depends on five independent kinematic variables.  These can be chosen as $s$, $t_{pn}$, $|\vec{q}_{\text{l.f.}}|$, $\cos\theta$, and $\varphi$, as in reference Bromberg:1983he.  The scattering angles $\theta$ and $\varphi$ were introduced earlier in Figure 2, $|\vec{q}_{\text{l.f.}}|$ is the magnitude of the incoming beam momentum in the lab frame, while $s$ and $t_{pn}$ are defined as follows:
@@
Neglecting background contributions, to be discussed later, the process $\pi^{-}p\to\pi^{+}\pi^{-}n$ at low $|t_{pn}|$ has long been known to be approximated by effective one-pion exchange (1PE), as suggested early on in reference Chew:1958wd.  This approximation is motivated by the observation that the pion is the lightest strongly interacting particle.  That said, a competing mechanism associated with isobar formation can also contribute to the same final state at low $|t_{pn}|$ (see Figure 4(b)), obscuring an interpretation based on 1PE.  However, experimental data at incoming beam momenta between 1 and 5 GeV already indicate a 1PE-dominated region, with higher momenta expected to further suppress isobar contributions Martin:1976mb.
Assuming 1PE dominance, the doubly inclusive differential cross section for the process $\pi^{-}p\to\pi^{+}\pi^{-}n$ is given by Martin:1976mb:
@@
and $\sigma_{\pi\pi}^{\text{{\it off}}}$ is the off-shell $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$ cross section. [^6] Details on how off-shell effects can be included can be found, for example, on page 33 in reference Martin:1976mb.
@@
Although 1PE is expected to dominate the process $\pi^{-}p\to\pi^{+}\pi^{-}n$ at low $|t_{pn}|$, it can at best provide only a qualitative description.  In particular, as observed in reference Stampke:1982cf, the $\pi^{-}p\to\pi^{+}\pi^{-}n$ data from reference Bromberg:1983he show deviations from equation (51), including a dependence on the angle $\varphi$, a non-vanishing behaviour as $|t_{pn}|\to 0$, and a steeper fall-off at larger $|t_{pn}|$ (see also reference Martin:1976mb).  As emphasised in reference Stampke:1982cf, these discrepancies indicate the need to account for additional mechanisms beyond 1PE, such as heavier-meson exchanges, finite-size contributions, and absorption effects.
Among possible modifications to the 1PE model, called backgrounds in the literature, absorption effects are known to capture the most features of the $\pi^{-}p\to\pi^{+}\pi^{-}n$ data Martin:1976mb. For this reason, we will focus on a modification of the 1PE model designed to incorporate the effects of pion absorption in the incoming and outgoing channels.  While certainly not the only mechanism required for a complete description of the data, we expect these effects to provide a useful approximation beyond the simple 1PE model.
Specifically, we consider the Poor Man’s Absorption (PMA) model, designed to provide a simplified description of the absorption effects mentioned above, with the doubly inclusive differential cross section for the process $\pi^{-}p\to\pi^{+}\pi^{-}n$ given by Martin:1976mb:
@@
We now turn to the experimental data found in reference Bromberg:1983he, where the observed $\pi\pi$ decay distributions from the process $\pi^{-}p\to\pi^{+}\pi^{-}n$ were reported in terms of a spherical harmonic expansion.  For $|t_{pn}|<0.15$ GeV2, only the $m=0$ and $1$ moments were significant, allowing the $\pi\pi$ decay distribution at fixed $s$, $t_{pn}$, and $|\vec{q}_{\text{l.f.}}|$ to be approximated by:
@@
where $I_{\text{P}}$ is proportional to the doubly inclusive differential cross section for the process $\pi^{-}p\to\pi^{+}\pi^{-}n$, and $I_{0}$ and $I_{1}$ denote the $m=0$ and $1$ moments, respectively.
In reference Bromberg:1983he, data for $I_{0}$ were presented, while $I_{1}$ data were not always available. The $I_{0}$ data were obtained from fits performed on $s^{1/2}$ and $\cos\theta$ bins over the interval $|t_{pn}|<0.15$ GeV2 (rather than at a single $|t_{pn}|$ value).  Measurements were taken at both incoming beam momenta of $|\vec{q}_{\text{l.f.}}|\approx 100$ GeV and $175$ GeV, enabling a simple qualitative check of whether the studied process $\pi^{-}p\to\pi^{+}\pi^{-}n$ is dominated by a $|\vec{q}_{\text{l.f.}}|$-independent $\pi\pi$ production mechanism, successfully carried out in Appendix B.  Finally, by comparing the reported $I_{0}$ data with the 1PE or PMA model, we can obtain indirect information on the scattering process $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$, as will be discussed in the next subsection.
@@ 4 Comparisons and Predictions
This section presents our model’s predictions using the bottum-up hard-wall AdS model for high-energy pion scattering. In Section 4.1, we compare our predictions for the angular dependence of $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$ scattering with experimental data extracted from $\pi^{-}p\to\pi^{+}\pi^{-}n$ measurements reported in reference Bromberg:1983he. In Section 4.2, we present predictions for all 22-to-22 pion scattering processes in one place.
@@ 4.1 Comparisons to Data
As outlined in Section 3, the available data come from $\pi^{-}p\to\pi^{+}\pi^{-}n$ measurements and can be related to $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$ scattering only with additional models.  In our analysis, we use both of the models introduced earlier.  From the discussion around equation (56), we see that the available data consist of the components $I_{0}(\theta)$ and $I_{1}(\theta)$, which together form the $\pi\pi$ decay distribution equal to the doubly inclusive differential cross section for the process $\pi^{-}p\to\pi^{+}\pi^{-}n$ at fixed $s$, $t_{pn}$, and $|\vec{q}_{\text{l.f.}}|$ up to an overall normalisation constant Martin:1976mb.
For comparisons using the 1PE model, only the $I_{0}(\theta)$ component can be considered, as the model does not account for $\varphi$ dependence, highlighting its inability to completely describe the process $\pi^{-}p\to\pi^{+}\pi^{-}n$.  In contrast, the PMA model provides predictions for both $I_{0}(\theta)$ and $I_{1}(\theta)$.  However, since the available data in reference Bromberg:1983he include only $I_{0}(\theta)$, our comparison with the PMA model will also be limited to this component.
@@
$I_{0}(s,\theta)=\mathcal{C}_{0}\,\big|\mathcal{M}(s,\theta)\big|^{2},$ (57)
@@
$I_{0}(s,\theta)=\mathcal{C}_{1}\,\Bigg[\big|\mathcal{M}(s,\theta)\big|^{2}+\frac{\mathcal{C}_{2}}{\tilde{\alpha}^{\prime}s}\bigg|\frac{\partial\mathcal{M}(s,\theta)}{\partial\theta}\bigg|^{2}\Bigg],$ (58)
@@
which is the slope of the $\rho$-meson linear Regge trajectory Collins:1977jy; Ebert:2009ub; Sonnenschein:2018fph.
In Figure 5, we plot the fitting results for the 175 GeV data set. We can see that the best fits are, as expected, at the highest energies available, $\sqrt{s}=3.15$ and $3.45$ GeV, where ${\tilde{\alpha}}^{\prime}s\approx 9$–10 is beginning to be large. These last two fit remarkably well at a qualitative level, which is encouraging despite the many approximations involved. The high energy fits are better also in the sense that the errors in the fitting parameters are smaller, a sign that the fitting algorithm converges better with the given model. In Tables 1 and 2, we list the fitting parameters we obtained. Their precise values have no particular physical significance for our purposes, and we list them only for completeness. Additional plots, including those for the 100 GeV data set are in Appendix A.
@@
The Minkowski space string amplitude predicts a large number of zeros and hence dips at $\tilde{\alpha}^{\prime}u=-n$ for any integer $n$. In the older literature, this feature, known as the Veneziano dip, was observed in various scattering experiments, though not always for all values of $n$ Odorico:1971dd; Odorico:1972vh; Pennington:1973dz; Eguchi:1973iv; Stampke:1982cf. It was argued to be a consequence of the Veneziano model, whose pole structure also implies an analogous structure of zeros responsible for the dips. Even though the basis for this phenomenon is the high energy behaviour, it persists at low energies, as the analytic structure of the amplitude remains constrained.
On the other hand, the dip at $\tilde{\alpha}^{\prime}t=-1$ is a different feature from the Veneziano dips at constant $\tilde{\alpha}^{\prime}u$. Our holographic model, after the integration over the flat space amplitude that has many dips, has only one or two zeros, though remarkably the observed dip at $\tilde{\alpha}^{\prime}t\approx-1$ is retained. Its location also serves as an independent verification of the value we set for $\tilde{\alpha}^{\prime}$, as increasing or decreasing $\tilde{\alpha}^{\prime}$ will shift the dip to smaller or larger angles, respectively. The value ${\tilde{\alpha}}^{\prime}=0.9$ GeV-2 is consistent with the position of the dip in the data.
@@
An important consideration that we have to make when fitting the data is the mismatch in the Regge regime, causing a discrepancy with our model at low $t$, with the threshold being around $|t|/s\approx 0.1$. This will require us to to adjust our fitting strategy accordingly. In light of this, our preferred fitting strategy was to fit all points to the right of the dip, and one or two points immediately to the left of it, even when those were below the small angle threshold of $|t|/s\approx 0.1$.
@@
Figure 5: Fits for the 175 GeV dataset. The solid black line represents the best fit using the PMA model, while the grey band corresponds to one standard deviation of the fitting parameters in the PMA model. The dashed black like corresponds to a best fit using 1PE model. Data points shown in red were excluded from the fit. The experimental data were digitised from Figure 29 in reference Bromberg:1983he; it is worth noting that we were not able to recover all data points or errors faithfully, although such cases were rare. The $s$-values indicated above correspond to the binned $s$-values included in the experimental measurements, whereas our fits are made only to the average of these binned values.

### v2

@@ 2.3 Precise Formulation of Scattering Amplitudes
In the strict high-energy fixed-angle limit $y_{0}\to\infty$, our predictions do not depend on the exact value of $\tilde{\alpha}^{\prime}$, as long as it remains finite. Hence, the leading behavior does not depend on the value of the IR cut-off, or more generally, on how confinement is realized in the model.
When examining pion scattering, we must also account for the pion’s isospin (realized in our holographic model using Chan–Paton factors, which we omitted earlier in Section 2.2). To do this, we follow the standard procedure Ecker (1995); Igi and Hikasa (1999), in which amplitudes of various pion scattering processes are represented by a single analytic function, the invariant amplitude:
@@ 2.3.1 Regularisation and Numerical Methods
As mentioned, evaluating equation (36) at integer values of $\tilde{\alpha}^{\prime}s$ leads to singularities that cannot be regularised.  These can be shown to correspond to logarithmic $s$-channel singularities Bianchi et al. (2022), although we do not expect that these singularities are physical, as they are restricted to finite $s$.  On the other hand, as long as $s$ does not coincide with a singularity, i.e.  when $\tilde{\alpha}^{\prime}s\neq n$, all other singularities can be regularised to give a well-defined integral, from which the angular dependence of the amplitude can be computed.  One way to achieve this is by shifting all the poles by $i\epsilon$, effectively giving a small width to all the mesons on the Regge trajectory. [^5] Phenomenologically, we expect finite widths, since all of the states on the $\rho$-meson Regge trajectory are unstable and decay.  In fact, the string decay width is expected to increase with spin, being proportional to the string length, which scales linearly with the mass Dai and Polchinski (1989); Peeters et al. (2006); Iengo and Russo (2006); Sonnenschein and Weissman (2018).  In the ’t Hooft limit, the excited states are stable Witten (1979), since the width is $\sim 1/N$, and the width functions serve only as regulators, which are eventually taken to zero.  A finite physical limit can then be defined by taking the limit $\epsilon\to 0$ in an appropriate manner.
@@
Now, since our predictions are based on tree-level superstring scattering amplitudes, the pion scattering amplitude $\mathcal{A}$ must be real.  This follows from the unitarity constraint Peskin and Schroeder (1995):
@@
where $T$ is the transfer matrix related to $\mathcal{A}$ via the LSZ formula, so $\mathcal{A}\sim T$.  At tree level, the right-hand side above vanishes, implying that unitarity forces $\mathcal{A}$ to be real. This connection between unitarity and the reality of the amplitude can also be seen from the fact that, at large-NN, the poles of the amplitude must lie on the positive real axis Caron-Huot et al. (2017).
@@ 3 Background Material on Experimental Data
In this section, we describe how the data for $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$ scattering can be extracted, under certain approximations, from the process $\pi^{-}p\to\pi^{+}\pi^{-}n$ studied in reference Bromberg and others (1984a). To this end, we introduce two models of increasing generality for extracting the data that, together with our holographic model, will form the basis for the analysis in Section 4.
In the experiment reported in reference Bromberg and others (1984a), the process $\pi^{-}p\to\pi^{+}\pi^{-}n$ was studied at incident beam momenta $|\vec{q}_{\text{l.f.}}|$ of approximately 100 GeV and 175 GeV in the lab frame. The incident beam was primarily composed of pions (approximately 90% at 100 GeV and 92% at 175 GeV), while liquid hydrogen was used as the target.  All the reported data were corrected for known experimental biases and acceptance losses, with uncertainties in these corrections remaining significantly smaller than the statistical errors across all data sets. Following the conventions of reference Bromberg and others (1984a), we frequently use the Gottfried–Jackson frame in this section (see Figure 2).  Additional information about the experiment can be found in references Fredericksen (1982); Stampke (1982); Bromberg and others (1984b, c), while related studies are reported in references Aderholz and others (1964b, a); Bondar and others (1966); Hyams et al. (1968); Robertson et al. (1973); Grayer and others (1974); Corden and others (1978); Wicklund et al. (1978); Weilhammer and others (1979).
@@
After averaging over the target proton spin and summing over the recoil neutron spin, the process $\pi^{-}p\to\pi^{+}\pi^{-}n$ depends on five independent kinematic variables.  These can be chosen as $s$, $t_{pn}$, $|\vec{q}_{\text{l.f.}}|$, $\cos\theta$, and $\varphi$, as in reference Bromberg and others (1984a).  The scattering angles $\theta$ and $\varphi$ were introduced earlier in Figure 2, $|\vec{q}_{\text{l.f.}}|$ is the magnitude of the incoming beam momentum in the lab frame, while $s$ and $t_{pn}$ are defined as follows:
@@
Neglecting background contributions, to be discussed later, the process $\pi^{-}p\to\pi^{+}\pi^{-}n$ at low $|t_{pn}|$ has long been known to be approximated by effective one-pion exchange (1PE), as suggested early on in reference Chew and Low (1959).  This approximation is motivated by the observation that the pion is the lightest strongly interacting particle.  That said, a competing mechanism associated with isobar formation can also contribute to the same final state at low $|t_{pn}|$ (see Figure 4(b)), obscuring an interpretation based on 1PE.  However, experimental data at incoming beam momenta between 1 and 5 GeV already indicate a 1PE-dominated region, with higher momenta expected to further suppress isobar contributions Martin et al. (1976).
Assuming 1PE dominance, the doubly inclusive differential cross section for the process $\pi^{-}p\to\pi^{+}\pi^{-}n$ is given by Martin et al. (1976):
@@
and $\sigma_{\pi\pi}^{\text{{\it off}}}$ is the off-shell $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$ cross section. [^6] Details on how off-shell effects can be included can be found, for example, on page 33 in reference Martin et al. (1976).
@@
Although 1PE is expected to dominate the process $\pi^{-}p\to\pi^{+}\pi^{-}n$ at low $|t_{pn}|$, it can at best provide only a qualitative description.  In particular, as observed in reference Stampke (1982), the $\pi^{-}p\to\pi^{+}\pi^{-}n$ data from reference Bromberg and others (1984a) show deviations from equation (51), including a dependence on the angle $\varphi$, a non-vanishing behaviour as $|t_{pn}|\to 0$, and a steeper fall-off at larger $|t_{pn}|$ (see also reference Martin et al. (1976)).  As emphasised in reference Stampke (1982), these discrepancies indicate the need to account for additional mechanisms beyond 1PE, such as heavier-meson exchanges, finite-size contributions, and absorption effects.
Among possible modifications to the 1PE model, called backgrounds in the literature, absorption effects are known to capture the most features of the $\pi^{-}p\to\pi^{+}\pi^{-}n$ data Martin et al. (1976). For this reason, we will focus on a modification of the 1PE model designed to incorporate the effects of pion absorption in the incoming and outgoing channels.  While certainly not the only mechanism required for a complete description of the data, we expect these effects to provide a useful approximation beyond the simple 1PE model.
Specifically, we consider the Poor Man’s Absorption (PMA) model, designed to provide a simplified description of the absorption effects mentioned above, with the doubly inclusive differential cross section for the process $\pi^{-}p\to\pi^{+}\pi^{-}n$ given by Martin et al. (1976):
@@
We now turn to the experimental data found in reference Bromberg and others (1984a), where the observed $\pi\pi$ decay distributions from the process $\pi^{-}p\to\pi^{+}\pi^{-}n$ were reported in terms of a spherical harmonic expansion.  For $|t_{pn}|<0.15$ GeV2, only the $m=0$ and 11 moments were significant, allowing the $\pi\pi$ decay distribution at fixed $s$, $t_{pn}$, and $|\vec{q}_{\text{l.f.}}|$ to be approximated by:
@@
where $I_{\text{P}}$ is proportional to the doubly inclusive differential cross section for the process $\pi^{-}p\to\pi^{+}\pi^{-}n$, and $I_{0}$ and $I_{1}$ denote the $m=0$ and 11 moments, respectively.
In reference Bromberg and others (1984a), data for $I_{0}$ were presented, while $I_{1}$ data were not always available. The $I_{0}$ data were obtained from fits performed on $s^{1/2}$ and $\cos\theta$ bins over the interval $|t_{pn}|<0.15$ GeV2 (rather than at a single $|t_{pn}|$ value).  Measurements were taken at both incoming beam momenta of $|\vec{q}_{\text{l.f.}}|\approx 100$ GeV and $175$ GeV, enabling a simple qualitative check of whether the studied process $\pi^{-}p\to\pi^{+}\pi^{-}n$ is dominated by a $|\vec{q}_{\text{l.f.}}|$-independent $\pi\pi$ production mechanism, successfully carried out in Appendix B.  Finally, by comparing the reported $I_{0}$ data with the 1PE or PMA model, we can obtain indirect information on the scattering process $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$, as will be discussed in the next subsection.
@@ 4 Comparisons and Predictions
This section presents our model’s predictions using the bottom-up hard-wall AdS model for high-energy pion scattering. In Section 4.1, we compare our predictions for the angular dependence of $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$ scattering with experimental data extracted from $\pi^{-}p\to\pi^{+}\pi^{-}n$ measurements reported in reference Bromberg and others (1984a). In Section 4.2, we present predictions for all 22-to-22 pion scattering processes in one place.
@@ 4.1 Comparisons to Data
As outlined in Section 3, the available data come from $\pi^{-}p\to\pi^{+}\pi^{-}n$ measurements and can be related to $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$ scattering only with additional models.  In our analysis, we use both of the models introduced earlier.  From the discussion around equation (56), we see that the available data consist of the components $I_{0}(\theta)$ and $I_{1}(\theta)$, which together form the $\pi\pi$ decay distribution equal to the doubly inclusive differential cross section for the process $\pi^{-}p\to\pi^{+}\pi^{-}n$ at fixed $s$, $t_{pn}$, and $|\vec{q}_{\text{l.f.}}|$ up to an overall normalisation constant Martin et al. (1976).
For comparisons using the 1PE model, only the $I_{0}(\theta)$ component can be considered, as the model does not account for $\varphi$ dependence, highlighting its inability to completely describe the process $\pi^{-}p\to\pi^{+}\pi^{-}n$.  In contrast, the PMA model provides predictions for both $I_{0}(\theta)$ and $I_{1}(\theta)$.  However, since the available data in reference Bromberg and others (1984a) include only $I_{0}(\theta)$, our comparison with the PMA model will also be limited to this component.
@@
$I_{0}(s,\theta)=\mathcal{C}_{0}\,\big|\mathcal{M}(s,\theta)\big|^{2}\,.$ (57)
@@
$I_{0}(s,\theta)=\mathcal{C}_{1}\,\Bigg[\big|\mathcal{M}(s,\theta)\big|^{2}+\frac{\mathcal{C}_{2}}{\tilde{\alpha}^{\prime}s}\bigg|\frac{\partial\mathcal{M}(s,\theta)}{\partial\theta}\bigg|^{2}\Bigg]\,.$ (58)
@@
which is the slope of the $\rho$-meson linear Regge trajectory Collins (1977); Ebert et al. (2009); Sonnenschein and Weissman (2019).
In Figure 5, we plot the fitting results for the 175 GeV data set. We can see that the best fits are, as expected, at the highest energies available, $\sqrt{s}=3.15$ and $3.45$ GeV, where ${\tilde{\alpha}}^{\prime}s\approx 9$–10 is beginning to be large. These last two fit remarkably well at a qualitative level, which is encouraging despite the many approximations involved. The high-energy fits are better also in the sense that the errors in the fitting parameters are smaller, a sign that the fitting algorithm converges better with the given model. In Tables 1 and 2, we list the fitting parameters we obtained. Their precise values have no particular physical significance for our purposes, and we list them only for completeness. Additional plots, including those for the 100 GeV data set are in Appendix A.
@@
The Minkowski space string amplitude predicts a large number of zeros and hence dips at $\tilde{\alpha}^{\prime}u=-n$ for any integer $n$. In the older literature, this feature, known as the Veneziano dip, was observed in various scattering experiments, though not always for all values of $n$ Odorico (1971, 1972); Pennington (1973); Eguchi et al. (1974); Stampke (1982). It was argued to be a consequence of the Veneziano model, whose pole structure also implies an analogous structure of zeros responsible for the dips. Even though the basis for this phenomenon is the high-energy behaviour, it persists at low energies, as the analytic structure of the amplitude remains constrained.
On the other hand, the dip at $\tilde{\alpha}^{\prime}t\approx-1$ is a different feature from the Veneziano dips at constant $\tilde{\alpha}^{\prime}u$. Our holographic model, after the integration over the flat space amplitude that has many dips, has only one or two zeros, though remarkably the observed dip at $\tilde{\alpha}^{\prime}t\approx-1$ is retained. Its location also serves as an independent verification of the value we set for $\tilde{\alpha}^{\prime}$, as increasing or decreasing $\tilde{\alpha}^{\prime}$ will shift the dip to smaller or larger angles, respectively. The value ${\tilde{\alpha}}^{\prime}=0.9$ GeV-2 is consistent with the position of the dip in the data.
@@
An important consideration that we have to make when fitting the data is the mismatch in the Regge regime, causing a discrepancy with our model at low $t$, with the threshold being around $|t|/s\approx 0.1$. This will require us to adjust our fitting strategy accordingly. In light of this, our preferred fitting strategy was to fit all points to the right of the dip, and one or two points immediately to the left of it, even when those were below the small angle threshold of $|t|/s\approx 0.1$.
@@
Figure 5: Fits for the 175 GeV dataset. The solid black line represents the best fit using the PMA model, while the grey band corresponds to one standard deviation of the fitting parameters in the PMA model. The dashed black line corresponds to a best fit using the 1PE model. Data points shown in red were excluded from the fit. The experimental data were digitised from Figure 29 in reference Bromberg and others (1984a); it is worth noting that we were not able to recover all data points or errors faithfully, although such cases were rare. The $s$-values indicated above correspond to the binned $s$-values included in the experimental measurements, whereas our fits are made only to the average of these binned values.